# Quark-diquark effective mass formalism for heavy baryon spectroscopy

### 01 Mar 2026 v1

@@ Abstract
We present a comprehensive study of heavy flavor baryons within a quark-diquark effective mass formalism, formulated using the effective masses of quarks and diquarks inside baryons. We predict the masses of $J^{P}=\frac{1}{2}^{+}$ and $J^{P}=\frac{3}{2}^{+}$ states under two complementary scenarios: the quark-quark interaction picture (Scenario I) and the quark-diquark interaction picture (Scenario II). Scenario I considers all possible quark-quark correlations, while Scenario II employs fixed spin-flavor diquark configurations, treating the baryon as an effective quark-diquark system. Using current experimental data, we estimate constituent quark masses, diquark masses, and hyperfine interactions, and include a mass-dependent binding energy term to account for short-range chromoelectric effects. The analysis shows that the binding energy plays a crucial role in describing heavy-heavy diquarks. The predicted baryon masses show excellent agreement with experimental and lattice QCD results across the charm and bottom sectors.
@@ I Introduction
Recent years have seen remarkable progress in the spectroscopy of heavy flavor baryons, largely driven by LHCb and CMS. Nearly the full singly heavy baryon spectrum has now been measured, except for $\Sigma_{b}^{(*)0}$, $\Xi_{b}^{\prime 0}$, and $\Omega_{b}^{*-}$ [1]. A major milestone was the observation of $\Xi_{cc}^{++}$ and its isospin partner $\Xi_{cc}^{+}$ [2, 3, 4, 5]. Searches for the doubly heavy baryons $\Xi_{cb}^{+(0)}$, $\Omega_{cb}^{0}$, and $\Omega_{cc}^{+}$ are ongoing [6, 7, 8, 9]. The LHCb Upgrade II, with higher luminosity, is expected to reveal additional states in the HL-LHC era [10], while future $e^{+}e^{-}$ colliders promise detailed studies of doubly and triply heavy baryons [11, 12, 13, 14, 15]. The discovery of these heavy baryons provides crucial insight into heavy-quark dynamics and QCD confinement, bridging conventional baryons and exotic multiquark systems.
@@
The observation of the open-flavor doubly heavy tetraquark $T_{cc}(3875)^{+}$ with a precisely measured mass of $3874.74(10)$ MeV [16], together with a growing set of tetraquark candidates [17, 18, 19], has strengthened the empirical link between conventional hadron spectroscopy and multiquark dynamics. More broadly, a wide range of manifestly exotic states has now been reported. These include charged hidden-flavor tetraquarks, such as $T_{c\bar{c}1}(3900)$ and their bottom counterparts, as well as fully heavy structures, including the di-$J/\psi$ resonances $X(6600)$ and $X(6900)$ observed by CMS, LHCb, and ATLAS [20, 21, 22]. The emerging hidden-charm pentaquark spectrum [23, 24, 25, 26, 27, 28] further reinforces the view that multiquark hadrons are a persistent feature of the QCD spectrum. These observations pose a pressing theoretical question: can the same quark correlations that govern conventional baryon structure also organize the spectroscopy of tetraquark and pentaquark states? Answering this question requires a framework in which the building blocks of exotic hadrons, namely diquarks, are not treated as free parameters but are instead calibrated against the precision baryon data that are now available.
[…]
@@
Within this context, diquark correlations offer a simplified yet quantitatively testable description of baryon structure, serving as a bridge between conventional and exotic hadrons in the QCD landscape. In this work, we develop a quark-diquark effective mass formalism (QDEMF) and apply it in two complementary scenarios. Scenario I treats all possible diquark combinations arising from quark-quark interactions, offering a comprehensive (broader) dynamical picture. Scenario II fixes the diquark spin-flavor content, furnishing a refined and physically transparent quark-diquark interaction picture. We introduce a mass-dependent binding energy term to capture short-range chromoelectric effects and the transition from chromomagnetic to color-Coulomb dominance in accordance with heavy-quark spin symmetry (HQSS). Our goal is to establish a quantitatively controlled baryon spectroscopy framework in which diquark properties, masses, spin assignments, and color representations can be extracted directly from data with minimal dynamical assumptions and subsequently deployed as calibrated, parameter-free inputs for multiquark spectroscopy.
The novel elements of this study are multifold: (i) prediction of low-lying heavy-baryon masses across singly, doubly, and triply heavy sectors within a single, parameter-light framework; (ii) the first systematic extraction and analysis of effective diquark masses and flavor content in both Scenario I and Scenario II; and (iii) an explicit, unified treatment of axial-vector versus scalar diquarks and a transparent description of hyperfine (spin-spin) splittings, demonstrating the expected $1/(m_{i}m_{j})$ suppression and the approach to HQSS in the heavy-quark limit. We demonstrate that the baryon-calibrated diquark spectrum provides a self-consistent and symmetry-constrained foundation for baryon spectroscopy within the quark-diquark framework. The diquark masses and chromomagnetic couplings extracted from the baryon sector emerge as internally stable effective degrees of freedom, yielding a coherent description of spin-dependent baryon systematics without additional phenomenological inputs. This establishes the calibration as a structurally robust spectroscopic baseline rather than a sector-specific fit. [^1] The broader applicability of the calibrated diquark parameters to multiquark spectroscopy is explored in a companion study [57], where the extracted mass scales and hyperfine couplings are employed without refitting to construct tetraquark and pentaquark spectra across multiple flavor sectors. Within the present work, we therefore restrict attention to baryon spectroscopy, using the resulting consistency and symmetry patterns as the primary validation of the diquark-based description.
@@ II.1 Quark-diquark effective mass formalism
The antitriplet color state, with a coefficient of 2/3-2/3, indicates an attractive interaction where two quarks are bound to form a stable configuration. Conversely, gluon exchange is repulsive in the sextet state due to its +1/3+1/3 coefficient [59, 60, 61, 62]. Consequently, diquarks must be in a $\overline{3}_{c}$ state to form a color singlet baryon when combined with a $3_{c}$ quark. This allows for treating a baryon as a two-body system where a quark within the baryon perceives the diquark as an antiquark, behaving like a $\overline{3}_{c}$ state, analogous to the meson system [63].
@@
The baryon mass, though fundamentally governed by the QCD Hamiltonian, is often estimated using a non-relativistic three-quark model. In this approach, a unitary transformation projects the QCD Hamiltonian onto a parametrized mass operator acting on quark-level spin-flavor wave functions [68, 69]. This operator effectively encodes key QCD dynamics, including confinement, chiral symmetry breaking, gluon exchange, and diquark correlations. Within this framework, the spin-flavor structure dictated by $SU(6)$ symmetry determines the relative contributions of scalar and axial-vector diquarks, weighted by CG coefficients. These coefficients influence the operator’s expectation values, thereby affecting the baryon mass and hyperfine splittings [66]. Importantly, all three diquark pairings contribute via quark-quark interactions mediated by one-gluon exchange (OGE).
We introduce QDEMF as a refined framework for estimating baryon masses by incorporating diquark correlations directly into the effective mass scheme (EMS) [70, 71, 72, 73]. In this framework, various interaction contributions to baryon masses are absorbed through renormalization of constituent quark masses. In EMS, the constituent masses thus incorporate all spin-independent effects of confinement and color-Coulomb dynamics, while treating strong hyperfine interactions explicitly via spin-spin correlations. This approach models baryons as three constituent quarks interacting via spin-dependent OGE,
@@
where $b_{ij}$ represents the strong hyperfine interaction between quarks $i$ and $j$ with a strength given by Eq. (2). These mass expressions include spin-dependent corrections arising from OGE, consistent with the QCD-inspired effective framework. Unlike Scenario I, the diquark masses of $D^{\{ij\}}$ and $D^{[ij]}$ in Scenario II exclude the multiplicative factor $3$ in the hyperfine term $b_{ij}$, avoiding the full absorption of hyperfine contributions into the diquark. This separation enables the residual interaction between the diquark and the third quark to be treated explicitly at the effective quark-diquark level. The hyperfine coupling between a diquark $D$ (composed of quarks $i$ and $j$) and the third quark $k$ is denoted $b_{\{ij\}k}$, with corresponding spin operators $\bm{s}_{D}$ and $\bm{s}_{k}$. As indicated by the color factor $\langle\vec{\lambda}_{\text{diquark}}\cdot\vec{\lambda}_{k}\rangle_{\text{singlet}}$ in Eq. (6), this interaction mirrors the color structure of the quark-antiquark coupling in mesons, allowing $b_{\{ij\}k}$ to be modeled analogously [79, 80, 81, 82, 83, 73]. In this construction, the underlying short-distance chromomagnetic scale is inherited from the same OGE interaction that governs quark-quark hyperfine splittings, with the contact matrix element $\langle\psi|\delta^{3}(\vec{r})|\psi\rangle$ probing only the resolved short-distance limit. The resulting coefficient therefore encodes the same short-distance chromomagnetic interaction in quark-diquark degrees of freedom, without introducing an additional dynamical scale. Accordingly,
$b_{\{ij\}k}=\frac{32\pi\alpha_{s}}{9m_{\{ij\}}m_{k}}\langle\psi|\delta^{3}(\vec{r})|\psi\rangle,$ (13)
and therefore, the general expressions for baryon masses can be written in terms of the effective masses of the diquark and quark constituents within the baryon $B(ijk)$, as follows:

### 09 Apr 2026 v2

@@ Abstract
We develop a quark–diquark effective mass formalism for heavy-flavor baryon spectroscopy and apply it to the $J^{P}=\tfrac{1}{2}^{+}$ and $J^{P}=\tfrac{3}{2}^{+}$ spectra across the singly, doubly, and triply heavy sectors. The analysis is carried out in two complementary scenarios: Scenario I treats all quark–quark diquark channels dynamically, while Scenario II restricts the dynamics to scalar and axial-vector diquarks, providing a more selective and physically transparent description. Constituent quark masses, effective diquark masses, and chromomagnetic couplings are extracted from known heavy-baryon masses, with the couplings determined solely by the quark content of each state and no sector-dependent adjustment introduced. A mass-dependent binding term is implemented to account for spin-independent chromoelectric effects and to describe the transition from chromomagnetic to color-Coulomb dominance across the light-to-heavy quark regime, ensuring consistency with heavy-quark spin symmetry in the heavy-quark limit. The resulting predictions are in good agreement with available experimental measurements and lattice QCD results across both charm and bottom sectors. The extracted diquark parameters remain stable across all heavy-flavor sectors, establishing the present framework as a symmetry-constrained spectroscopic baseline for heavy-baryon structure.
@@ I Introduction
Recent years have seen remarkable progress in the spectroscopy of heavy flavor baryons, largely driven by LHCb and CMS. Nearly the full ground-state singly heavy baryon spectrum has now been measured [91]. A major milestone was the observation of $\Xi_{cc}^{++}$ and its isospin partner $\Xi_{cc}^{+}$ [5, 8, 19, 12, 9]. Searches for the doubly heavy baryons $\Xi_{cb}^{+(0)}$, $\Omega_{cb}^{0}$, and $\Omega_{cc}^{+}$ are ongoing [17, 14, 13, 10]. The LHCb Upgrade II, with higher luminosity, is expected to reveal additional states in the HL-LHC era [103], while future $e^{+}e^{-}$ colliders promise detailed studies of doubly and triply heavy baryons [108, 21, 36, 47, 20]. The discovery of these heavy baryons provides crucial insight into heavy-quark dynamics and quantum chromodynamics (QCD) confinement, bridging conventional and exotic multiquark systems.
@@
A pressing theoretical question now confronts this experimental richness: can the same quark correlations that govern conventional baryon structure also organize the spectroscopy of tetraquark and pentaquark states? The observation of the open-flavor doubly heavy tetraquark $T_{cc}(3875)^{+}$ with a precisely measured mass of $3874.74(10)$ MeV [16], together with a growing set of tetraquark candidates [18, 2, 67], has strengthened the empirical link between conventional hadron spectroscopy and multiquark dynamics. More broadly, a wide range of manifestly exotic states has now been reported. These include charged hidden-flavor tetraquarks, such as $Z_{c}(3900)$ and their bottom counterparts, as well as fully heavy resonant structures, including the di-$J/\psi$ resonances $X(6600),X(6900)$, and $X(7100)$ observed by CMS, LHCb, and ATLAS [62, 7, 1]. The emerging hidden-charm pentaquark spectrum [22, 3, 15, 11, 6, 4] further reinforces the view that multiquark hadrons are a persistent feature of the QCD spectrum. These observations have motivated numerous theoretical approaches, including potential models, relativistic quark models, QCD sum rules, chiral quark models, and lattice QCD [60, 107, 58, 87, 72, 96, 59, 54, 52, 68, 106, 99, 100, 98, 97, 109, 101, 28, 26, 27, 79, 92, 94, 31, 32, 86, 83, 82]. While these approaches have yielded valuable sector-specific insights, none provides a single framework in which diquark properties are extracted from and validated against the current baryon spectrum and subsequently applied to multiquark systems. Establishing such a calibration pipeline motivates the present work.
Within this context, diquark correlations [33, 64, 30] provide a robust, quantitatively testable description of baryon structure and serve as a natural bridge between conventional and exotic hadrons [80, 69, 55, 76, 40]. Recent CMS limits on quark substructure down to $10^{-20}$ m [61] preclude spatial smearing at TeV scales, justifying the contact-interaction limit in our chromomagnetic interaction and reinforcing the treatment of the diquark as a dynamically compact cluster. This interpretation is further supported by flavor-decomposition analyses of nucleon electromagnetic form factors at high momentum transfers [33, 35]. Additional evidence comes from heavy-baryon production mechanisms. For instance, double-charm production at LHCb proceeds predominantly through the formation of a compact heavy diquark that subsequently hadronizes [81, 102, 39, 78].
Recent analysis using the Gaussian expansion method have identified clear diquark correlations within baryons. In particular, the scalar-isoscalar diquark is found to have lower energy and a more compact spatial extent, consistent with the notion of a “good diquark” [110]. For QQQQ-$q$ systems, increasing the heavy-quark mass reduces the diquark separation and enhances the diquark clustering effect, supporting a compact heavy diquark picture. The quark-diquark framework has become standard for describing triply heavy baryons [59], constructing unified Regge trajectories for heavy-heavy systems [105, 41], and formulating chiral effective theories of diquarks [73]. Collectively, these developments show that the diquark configuration, by capturing dominant color-spin correlations and respecting HQSS, is a natural and necessary building block for extending QCD dynamics from conventional baryons to multiquark exotics.
@@
Thus, calibrating diquark properties directly from the baryon spectrum is a phenomenological necessity, not a mere modeling simplification. Motivated by this, we develop a quark-diquark effective mass formalism (QDEMF) in two complementary scenarios. Scenario I treats all possible diquark combinations arising from quark-quark interactions, providing a dynamically inclusive description. Scenario II instead fixes the diquark spin-flavor structure, restricting the dynamics to flavor-antisymmetric scalar and flavor-symmetric axial-vector diquarks, thereby furnishing a more selective and physically transparent interaction picture. To account for the changing dynamical regime between light and heavy quarks, we introduce a mass-dependent binding energy term [70]. This term captures short-range chromoelectric effects and the gradual transition from chromomagnetic dominance to spin-independent color-Coulomb dynamics in accordance with heavy-quark spin symmetry (HQSS).
The present framework yields several concrete outcomes. First, it provides predictions for the low-lying heavy-baryon spectrum across singly, doubly, and triply heavy sectors within a single parameter-light formulation. Second, it enables a systematic extraction and analysis of effective diquark masses and flavor content within the two complementary scenarios introduced above, with scalar and axial-vector contributions explicitly disentangled across all heavy-flavor sectors. Third, the framework provides a unified description of both scalar and axial-vector diquarks together with a transparent accounting of hyperfine (spin-spin) splittings. The characteristic $1/(m_{i}m_{j})$ chromomagnetic scaling arises naturally from the underlying one-gluon exchange (OGE) interaction and is reproduced consistently across all flavor sectors as a direct consequence of the dynamics. Consequently, the correct HQSS degeneracy structure emerges automatically in the heavy-quark limit, rather than being imposed as an external constraint.
The chromomagnetic couplings are determined entirely by the quark content of each state, with no sector-dependent adjustment introduced at any stage, so that the extracted diquark masses and hyperfine couplings act as stable predictive parameters across the full heavy-baryon spectrum. Within the present work we restrict attention to baryon spectroscopy, using the resulting symmetry patterns as the primary internal validation of the quark-diquark framework. [^1] The broader applicability of the calibrated diquark parameters to multiquark spectroscopy is explored in a companion study [84], where the extracted mass scales and hyperfine couplings are employed without refitting to construct tetraquark and pentaquark spectra across multiple flavor sectors.
@@ II.1 Quark-diquark effective mass formalism
The antitriplet color state, with a coefficient of 2/3-2/3, indicates an attractive interaction where two quarks are bound to form a stable configuration. Conversely, gluon exchange is repulsive in the sextet state due to its +1/3+1/3 coefficient [24, 25, 70, 64, 34]. Consequently, diquarks must be in a $\overline{3}_{c}$ state to form a color singlet baryon when combined with a $3_{c}$ quark. This allows for treating a baryon as a two-body system where a quark within the baryon perceives the diquark as an antiquark, behaving like a $\overline{3}_{c}$ state, analogous to the meson system [90].
@@
The baryon mass, though fundamentally governed by the QCD Hamiltonian, is often estimated using a non-relativistic three-quark model. In this approach, a unitary transformation projects the QCD Hamiltonian onto a parametrized mass operator acting on quark-level spin-flavor wave functions [89, 44]. This operator effectively encodes key QCD dynamics, including confinement, chiral symmetry breaking, gluon exchange, and diquark correlations. Within this framework, the spin-flavor structure dictated by $SU(6)$ symmetry determines the relative contributions of scalar and axial-vector diquarks, weighted by CG coefficients. These coefficients influence the operator’s expectation values, thereby affecting the baryon mass and hyperfine splittings [71]. Importantly, all three diquark pairings contribute via quark-quark interactions mediated by OGE.
We introduce QDEMF as a refined framework for estimating baryon masses by incorporating diquark correlations directly into the effective mass scheme (EMS) [85, 63, 43, 42]. In the standard EMS, all spin-independent effects of confinement and color-Coulomb dynamics are absorbed into renormalized constituent quark masses, while strong hyperfine interactions are treated explicitly via spin–spin correlations. In the present QDEMF, we retain this hyperfine structure but depart from the standard scheme by computing the spin-independent binding energy contribution explicitly and separately, rather than absorbing it into the constituent masses [88, 89]. This separation allows the color-Coulomb dynamics to be tracked across the light-to-heavy quark transition without conflating spin-independent and spin-dependent contributions at the level of the effective diquark mass. This approach models baryons as three constituent quarks interacting via spin-dependent OGE,
@@
where $b_{ij}$ represents the strong hyperfine interaction between quarks $i$ and $j$ with a strength given by Eq. (2). These mass expressions include spin-dependent corrections arising from OGE, consistent with the QCD-inspired effective framework. Unlike Scenario I, the diquark masses of $D^{\{ij\}}$ and $D^{[ij]}$ in Scenario II exclude the multiplicative factor $3$ in the hyperfine term $b_{ij}$, avoiding the full absorption of hyperfine contributions into the diquark. [^4] In the present construction, spin-independent color-Coulomb contributions are consistently included in the baseline OGE-based baryon mass framework [70]. Confinement-induced binding-energy effects are evaluated independently and added as explicit corrections. The effective diquark mass therefore reflects the internal short-distance structure fixed by baryon spectroscopy. This separation enables the residual chromomagnetic interaction between the diquark and the third quark to be treated explicitly at the effective quark-diquark level. The hyperfine coupling between a diquark $D$ (composed of quarks $i$ and $j$) and the third quark $k$ is denoted $b_{\{ij\}k}$, with corresponding spin operators $\bm{s}_{D}$ and $\bm{s}_{k}$. As indicated by the color factor $\langle\vec{\lambda}_{\text{diquark}}\cdot\vec{\lambda}_{k}\rangle_{\text{singlet}}$ in Eq. (6), this interaction mirrors the color structure of the quark-antiquark coupling in mesons, allowing $b_{\{ij\}k}$ to be modeled analogously [93, 37, 74, 75, 51]. Within this construction, the short-distance chromomagnetic scale is derived directly from the OGE interaction, the same fundamental dynamics governing quark-quark hyperfine splittings. The contact matrix element, $\langle\psi|\delta^{3}(\vec{r})|\psi\rangle$, isolates the resolved short-distance limit of the wavefunction [42, 88], leading to the hyperfine interaction coefficient:
$b_{\{ij\}k}=\frac{32\pi\alpha_{s}}{9m_{\{ij\}}m_{k}}\langle\psi|\delta^{3}(\vec{r})|\psi\rangle.$ (13)
This hyperfine interaction coefficient naturally preserves the $1/(m_{i}m_{j})$ inverse-mass hierarchy characteristic of OGE-driven interactions. By formulating this coupling strictly within quark-diquark degrees of freedom, we maintain the parameter-light nature of the framework without introducing an additional dynamical scale. Consequently, the general mass expressions for a baryon $B(ijk)$ can be parameterized through the effective masses of its constituent diquark and quark degrees of freedom as follows: